there are 6 people on a bus. There are 2 choices for the meal being served. what is the probability that all 6 people will choose the same meal assuming choices are made independently?

Answer by stanbon(75887)

there are 6 people on a bus. There are 2 choices for the meal being served. what is the probability that all 6 people will choose the same meal assuming choices are made independently?
---------------
2(1/2)^6 = 2(1/64) = 1/32
----
It's the same as all six people tossing a coin
and all of them getting heads or all getting tails.
======================================================
Cheers,
Stan H.